# Beyond Binary: Towards Fine-Grained LLM-Generated Text Detection via Role Recognition and Involvement Measurement

### v1

@@ Abstract.
The rapid development of large language models (LLMs), like ChatGPT, has resulted in the widespread presence of LLM-generated content on social media platforms, raising concerns about misinformation, data biases, and privacy violations, which can undermine trust in online discourse. While detecting LLM-generated content is crucial for mitigating these risks, current methods often focus on binary classification, failing to address the complexities of real-world scenarios like human-AI collaboration. To move beyond binary classification and address these challenges, we propose a new paradigm for detecting LLM-generated content. This approach introduces two novel tasks: LLM Role Recognition (LLM-RR), a multi-class classification task that identifies specific roles of LLM in content generation, and LLM Influence Measurement (LLM-IM), a regression task that quantifies the extent of LLM involvement in content creation. To support these tasks, we propose LLMDetect, a benchmark designed to evaluate detectors’ performance on these new tasks. LLMDetect includes the Hybrid News Detection Corpus (HNDC) for training detectors, as well as DetectEval, a comprehensive evaluation suite that considers five distinct cross-context variations and multi-intensity variations within the same LLM role. This allows for a thorough assessment of detectors’ generalization and robustness across diverse contexts. Our empirical validation of 10 baseline detection methods demonstrates that fine-tuned Pre-trained Language Model (PLM)-based models consistently outperform others on both tasks, while advanced LLMs face challenges in accurately detecting their own generated content. Our experimental results and analysis offer insights for developing more effective detection models for LLM-generated content. This research enhances the understanding of LLM-generated content and establishes a foundation for more nuanced detection methodologies.  [^1] Our code and dataset are available in https://github.com/ZihaoCheng123/LLMDetect.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. INTRODUCTION
— Peter Steiner
Recent advances in generative large language models (LLMs) (OpenAI, 2024; Jiang et al., 2023; Dubey et al., 2024; DeepSeek-AI et al., 2024), such as GPT-4 (OpenAI, 2024) and LLaMA (Dubey et al., 2024), alongside the increasing availability of tools like ChatGPT [^2] https://chatgpt.com/ and Copilot [^3] https://copilot.microsoft.com/, have significantly reshaped the landscape of social media and web platforms (Cheng, 2024). These technologies facilitate the automated creation of extensive content with human-like fluency (Partadiredja et al., 2020), making LLM-generated posts, articles, and comments widely accessible and rapidly disseminated. The proliferation of such content has profoundly expanded its reach and influence, transforming the dynamics of online discourse.
However, these advancements also introduce significant risks, both in terms of information accuracy and public trust. While LLM-generated content can match the fluency of professional writing, it inevitably contains hallucinations (Bang et al., 2023; Augenstein et al., 2024)—misleading information that appears credible but lacks factual accuracy. A report by NewsGuard [^4] https://www.newsguardtech.com/special-reports/ai-tracking-center/ identified over 1,050 unreliable LLM-generated news websites, further undermining the already fragile information ecosystem. The rapid spread of such content across social media heightens the risk of misinformation (Su et al., 2024; Puccetti et al., 2024), challenging the accuracy and credibility of digital information. Additionally, LLM-generated content often exhibits inherent biases (Fang et al., 2024) and can be misused for malicious purposes (Menczer et al., 2023; Vykopal et al., 2023), further complicating efforts to maintain information integrity. These risks contribute to the erosion of public trust in media. According to the 2024 Digital News Report (Behre et al., 2024), global trust in news media has fallen to 40%, and the rise of LLM-generated content threatens to further weaken this fragile trust. As distinguishing between human-written and LLM-generated content becomes critical for preserving information integrity (Brewster et al., 2023; Yang and Menczer, 2024), current detection methods, which are largely limited to binary classification (Wang et al., 2023; Tripto et al., 2023; Verma et al., 2024; Wang et al., 2024), fail to distinguish the complexity of LLM-generated content like mixed human-LLM input.
@@
In this paper, we propose a new paradigm for detecting LLM-generated content that moves beyond the limitations of binary classification by considering both the LLM’s role and level of involvement in content creation, as depicted in Figure 1. LLMs often play diverse roles in assisting human authors, and to capture this complexity, we introduce two novel tasks. The first task, LLM Role Recognition (LLM-RR), is a multi-class classification task that identifies the specific roles played by LLMs in content generation, distinguishing between stages such as drafting and refinement. The second task, LLM Influence Measurement (LLM-IM), is a regression task designed to quantify the LLM involvement ratio in content creation, offering a nuanced measure of AI influence on the final output.
@@
  • We propose a new detection paradigm that moves beyond binary classification, introducing two novel tasks: LLM Role Recognition (LLM-RR) and LLM Influence Measurement (LLM-IM).
@@ 2.1. Detection tasks
Figure 1. The detection framework toward fine-grained LLM-generated text detection through role recognition and involvement evaluation.
@@ 2.2. Detection methods
Current detection methods can be broadly categorized into three types based on the features they rely on (Ghosal et al., 2023): watermarking-based detection methods, statistical outlier detection methods, and fine-tuning classifiers. (i) The watermark-based methods require embedding the signals that are invisible to humans into the AI-generated text and then detecting them based on these invisible token-level secret markers (Kirchenbauer et al., 2023). However, this method not only requires pre-editing that is not applicable to open-source models (Puccetti et al., 2024) but also affects the quality of model generation due to the insertion of watermarks (Tang et al., 2024). (ii) Statistical outlier detection methods focus on distinguishing whether a text is written by GPT based on the human features contained in the text. They adopt features ranging from shallows (entropy (Lavergne et al., 2008; Gehrmann et al., 2019), n-gram frequencies (badaskar-etal-2008-identifying), and perplexity (Beresneva, 2016)) to deeps such as using the absolute rank (Gehrmann et al., 2019), the Log Likelihood Ratio Ranking (LRR) by complementing Log Rank (Su et al., 2023) and the model’s log probability in regions of negative curvature (DetectGPT) (Mitchell et al., 2023). (iii) Supervised fine-tuning classifiers, trained on annotated data (Jawahar et al., 2020; Bakhtin et al., 2019; Sadasivan et al., 2023), have shown effectiveness in detecting LLM-generated text across domains, such as news (Zellers et al., 2019; Kumarage et al., 2023), social media (e.g., Twitter) (Fagni et al., 2021), and academic papers (Yang et al., 2023). However, these classifiers often overfit to specific domains, leading to poor performance on out-of-distribution data (Uchendu et al., 2020; Chakraborty et al., 2023), and their capabilities significantly degrade when applied to unseen datasets from different domains (Li et al., 2024).
Thus, assessing the transferability of detection models is essential for their practical application across diverse datasets and domains. In our work, we evaluate detection methods—including zero-shot LLMs, supervised feature-based, and PLM-based models—on our novel tasks and dataset. Our empirical validation provides insights into improving detection models for LLM-generated content, particularly with respect to enhancing generalization and robustness in diverse, real-world scenarios.
@@ 3. METHODOLOGY
Figure 1 illustrates our proposed detection framework for fine-grained detection of LLM-generated text, encompassing two proposed novel detection tasks (§3.1), and the LLMDect Benchmark (HNDC & DetectEval) for training and evaluation detection models (§3.2).
@@ 3.1. Detection Paradigm Definition
The current task of detecting LLM-generated text primarily relies on binary classification, determining whether a text is LLM-generated or not. In this traditional binary LLM detection task, given a dataset $\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{N}$, where $x_{i}$ denotes the text content and $y_{i}\in\left\{0,1\right\}$ indicates whether the text is LLM-generated. However, this approach focuses only on identifying LLM-generated content and is unable to distinguish more complex scenarios. For example, in LLM-assisted writing, users may employ LLMs to refine or slightly modify sentence structures for improved fluency, which is different from cases where the LLM generates the entire text. To overcome these limitations, we propose two new detection tasks: LLM Role Recognition (LLM-RR), a multi-class classification task, and LLM Involvement Measurement (LLM-IM), a regression task.
@@ 3.1.1. LLM Role Recognition
LLM-RR aims to identify the specific role that LLM plays in text generation when used as an LLM-assisted writing tools. Unlike binary detection, where labels are binary, the label for each $x_{i}$ is defined as $y_{i}\in\left\{C_{1},C_{2},\dots,C_{k}\right\}$, indicating the specific role $C_{i}$ the LLM plays in generating text $x_{i}$. Examples of such roles include fully human-written text, LLM-generated content with minor human editing, human-led creation with LLM assistance, or fully LLM-generated text without human involvement, among others. The LLM-RR task can be defined as follows:
@@
where our objective is to find an optimal classifier $f(\cdot)$ that minimizes the overall average misclassification rate. The indicator function $\mathbf{1}\{f(x_{i})\neq y_{i}\}$ equals 1 when the classifier $f$ assigns an incorrect label to the input $x_{i}$, and 0 if the label is correct. The text $x_{i}$ follows a distinct distribution $\mathscr{F}_{k}$ conditioned on its category label $y_{i}$:
@@ 3.1.2. LLM Involvement Measurement
While LLM-RR provides greater granularity compared to binary detection, it also has limitations. First, user interactions with LLMs in real-world applications are complex, making it challenging to accurately define all possible LLM roles. Additionally, even within the same LLM role, the degree of the LLM’s contribution can vary, adding further complexity to detection. Therefore, we propose a new task, LLM Involvement Measurement (LLM-IM), to address these challenges. LLM-IM quantifies the degree of LLM involvement in the generated text and is framed as a regression task. In this task, the dataset $\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{N}$ features labels $y_{i}$ that represent a continuous value rather than discrete role categories. We define this value label as the LLM Involvement Ratio (LIR), a metric ranging from 0 to 1, where 0 indicates no LLM involvement, and 1 signifies that the text is entirely generated by the LLM. Specifically, it is calculated as follows:
@@
where our objective is to minimize the expected loss over the data distribution $\mathcal{D}$.
By integrating the LLM-RR and LLM-IM tasks, this framework offers a comprehensive and scalable approach to understanding both the roles and the extent of LLM involvement in content creation.
3.2. LLMDect Benchmark
To validate the effectiveness of our proposed detection paradigm, we construct LLMDect, a benchmark specifically designed to evaluate detection models across varying levels of LLM involvement. This benchmark encompasses four distinct roles in content creation: Human-Author, LLM-Creator, LLM-Polisher, and LLM-Extender. Each of the four roles represents a distinct level of LLM participation:
@@ 3.2. LLMDect Benchmark
Specifically, the LIR is defined as 0 for Human-Author text and 1 for LLM-Creator text, while for LLM-Polisher and LLM-Extender text, the LIR falls between 0 and 1. For LLM-Extender, the LIR value can be directly calculated using Equation 3. However, for LLM-Polisher, directly extracting $T_{LLM}$ is not feasible. Therefore, following the approach of Yang et al. (2023), we use the Jaccard distance to calculate the polish ratio, which serves as the LIR for this role.
In LLMDect, each Human-Author text is paired with three versions generated by the other LLM roles. Each text is annotated with its corresponding LLM role and associated LIR value. This dual annotation framework enables a comprehensive evaluation of both the role and extent of LLM involvement in content creation. The constructed LLMDetect benchmark comprises two key components: the Hybrid News Detection Corpus (HNDC), a diverse dataset designed for robust training and evaluation of detection methods, and DetectEval, a comprehensive evaluation suite featuring five distinct out-of-distribution settings and varying intensity levels within the same LLM role.
@@ 3.2.1. HNDC
b. LLM-Assisted News Generation
To generate LLM-generated news articles, we design distinct prompts based on the three proposed roles: (1) For LLM-Creator news, the prompt includes the title, available summary, topic category, and publication date to ensure factual reliability. (2) For LLM-Polisher news, the entire original article is provided. If the article is too long, it is segmented for polishing to avoid overly shortened outputs that may result from processing lengthy articles in one go. (3) For LLM-Extender news, we retain the first three sentences or up to one-third of the original text and instruct the LLM to generate the remaining content. To ensure high-quality generation, we employ role-playing prompts, assigning the LLM the role of a journalist. This approach leverages social role assignment, which has been shown to improve LLM performance consistently (Zheng et al., 2023).666 The prompts used for HNDC construction are provided in the Appendix A. We select LLaMa3 (Dubey et al., 2024) (Meta-Llama-3-8B-Instruct) as the writing assistant LLM.
@@ 3.2.2. DetectEval
DetectEval is a comprehensive evaluation suite designed to assess the transferability and robustness of detection models, specifically focusing on Cross-context variations and Multi-intensity Variations.
@@
Cross-context variations examine data diversity across five dimensions: content publication time, prompts for generation, source LLM, cultural differences, and content domain, resulting in five out-of-distribution settings: cross-time, cross-prompt, cross-source, cross-cultural, and cross-domain. The first four settings, similar to those in the HNDC, focus on LLM-generated content within the news domain, cross-prompt and cross-source settings are directly expanded based on the test dataset from HNDC.
Cross-time: HNDC’s pre-2019 articles reduce LLM involvement, but data leakage is still possible. To address this, we scrape 2024 New York Times articles and generate LLM content for different roles using the same method as HNDC.
Cross-prompt: LLM-assisted writing varies by prompt, even for the same role. We design five distinct prompts per LLM-assisted news role and pair them with news articles to create a diverse test set. Prompts are listed in Appendix B.
Cross-source: While HNDC initially used Llama-3-8B-Instruct, real-world scenarios involve stronger LLMs. We supplement HNDC test data with content generated from four more powerful models: Deepseek-v2, Meta-LLaMA-3-70B-Instruct, Claude-3.5-Sonnet, and GPT-4o.
Cross-cultural: Considering that writing and expression styles can vary across cultural contexts even within the same domain, we constructed a cross-cultural test set using news platforms (dai, 2017) from Germany, China, and Qatar. These countries reflect significant cultural diversity, especially in terms of language and values.
Cross-domain: Writing styles differ across domains. We construct a cross-domain dataset with text from thesis, story, and essay domains, sourced from CUDRT (Tao et al., 2024), MGTBench (He et al., 2023), and CDB (Liang et al., 2023). Specifically, we extract relevant text types from each dataset and supplement the missing categories using our defined methodology.
@@
Multi-intensity variations are introduced to address the fact that, even within the same LLM-assisted writing role, the level of LLM involvement can differ. We specifically construct test data with varying degrees of LLM involvement ratio for the LLM-extender and LLM-polisher roles.
Variable-Length Extension: For LLM-Extender role, we set up three truncation lengths to create variable-length extensions, allowing us to evaluate whether the detectors can identify the differing levels of content expansion. For the given text $x$, this process can be described as: $E^{\left(n\right)}(x),n\in\{Low,Medium,High\}$ [^7] We randomly retain part of an article’s initial sentences and ask LLMs to complete it. Low refers to retaining $[3,l/3]$ sentences, Medium retains $[l/3,2l/3]$, and High retains $[2l/3,l-3]$, where $l$ is the total number of sentences. , where $n$ denotes the truncation state of the text, and $E$ represents the process of text extension by LLMs.
Multi-Staged Polish: For LLM-polisher role, we apply a multi-staged polish process, iterating the text polishing up to six times to evaluate whether the detection methods can identify the increasing levels of refinement. For the given text $x$, this process can be described as: $P_{m}\left(x\right)$, where $m$ denotes the number of polish times, and $P$ represents the polishing process by LLMs.

### v2

@@ Abstract.
The rapid development of large language models (LLMs), like ChatGPT, has resulted in the widespread presence of LLM-generated content on social media platforms, raising concerns about misinformation, data biases, and privacy violations, which can undermine trust in online discourse. While detecting LLM-generated content is crucial for mitigating these risks, current methods often focus on binary classification, failing to address the complexities of real-world scenarios like human-LLM collaboration. To move beyond binary classification and address these challenges, we propose a new paradigm for detecting LLM-generated content. This approach introduces two novel tasks: LLM Role Recognition (LLM-RR), a multi-class classification task that identifies specific roles of an LLM in content generation, and LLM Involvement Measurement (LLM-IM), a regression task that quantifies the extent of LLM involvement in content creation. To support these tasks, we propose LLMDetect, a benchmark designed to evaluate detectors’ performance on these new tasks. LLMDetect includes the Hybrid News Detection Corpus (HNDC) for training detectors, as well as DetectEval, a comprehensive evaluation suite that considers five distinct cross-context variations and two multi-intensity variations within the same LLM role. This allows for a thorough assessment of detectors’ generalization and robustness across diverse contexts. Our empirical validation of 10 baseline detection methods demonstrates that fine-tuned Pre-trained Language Model (PLM)-based models consistently outperform others on both tasks, while advanced LLMs face challenges in accurately detecting their own generated content. Our experimental results and analysis offer insights for developing more effective detection models for LLM-generated content. This research enhances the understanding of LLM-generated content and establishes a foundation for more nuanced detection methodologies. [^1] Our code and dataset are available in https://github.com/ZihaoCheng123/LLMDetect.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714770isbn: 979-8-4007-1274-6/25/04ccs: Information systems Social mediaccs: Human-centered computing Social media
@@ 1. INTRODUCTION
Partadiredja et al. (2020)
Recent advances in generative large language models (LLMs) (OpenAI, 2024; Jiang et al., 2023; Dubey et al., 2024; DeepSeek-AI et al., 2024), such as GPT-4 (OpenAI, 2024) and LLaMA (Dubey et al., 2024), alongside the increasing availability of tools like ChatGPT and Copilot [^2] https://chatgpt.com/ and https://copilot.microsoft.com/, have significantly reshaped the landscape of social media and web platforms (Cheng, 2024). These technologies facilitate the automated creation of extensive content with human-like fluency (Partadiredja et al., 2020), making LLM-generated posts, articles, and comments widely accessible and rapidly disseminated. The proliferation of such content has profoundly expanded its reach and influence, transforming the dynamics of online discourse.
However, these advancements also introduce significant risks, both in terms of information accuracy and public trust. While LLM-generated content can match the fluency of professional writing, it inevitably contains hallucinations (Bang et al., 2023; Augenstein et al., 2024)—misleading information that appears credible but lacks factual accuracy. A report by NewsGuard [^3] https://www.newsguardtech.com/special-reports/ai-tracking-center/ identified over 1,050 unreliable LLM-generated news websites, further undermining the already fragile information ecosystem. The rapid spread of such content across social media heightens the risk of misinformation (Su et al., 2024; Puccetti et al., 2024a), challenging the accuracy and credibility of digital information. Additionally, LLM-generated content often exhibits inherent biases (Fang et al., 2024) and can be misused for malicious purposes (Menczer et al., 2023; Vykopal et al., 2023), further complicating efforts to maintain information integrity. These risks contribute to the erosion of public trust in media. According to the 2024 Digital News Report (Behre et al., 2024), global trust in news media has fallen to 40%, and the rise of LLM-generated content threatens to further weaken this fragile trust. As distinguishing between human-written and LLM-generated content becomes critical for preserving information integrity (Brewster et al., 2023; Yang and Menczer, 2024), current detection methods, which are largely limited to binary classification (Wang et al., 2023; Tripto et al., 2023; Verma et al., 2024; Wang et al., 2024; Cornelius et al., 2024; Dugan et al., 2024; Yu et al., 2024; Ma and Wang, 2024), fail to distinguish the complexity of LLM-generated content like mixed human-LLM input.
@@
In this paper, we propose a new paradigm for detecting LLM-generated content that moves beyond the limitations of binary classification by considering both the LLM’s role and level of involvement in content creation, as depicted in Figure 1. LLMs often play diverse roles in assisting human authors, and to capture this complexity, we introduce two novel tasks. The first task, LLM Role Recognition (LLM-RR), is a multi-class classification task that identifies the specific roles played by LLMs in content generation, distinguishing between stages such as drafting and refinement. The second task, LLM Involvement Measurement (LLM-IM), is a regression task designed to quantify the LLM involvement ratio in content creation, offering a nuanced measure of AI influence on the final output.
@@
  • We propose a new detection paradigm that moves beyond binary classification, introducing two novel tasks: LLM Role Recognition (LLM-RR) and LLM Involvement Measurement (LLM-IM).
@@ 2. RELATED WORK
Figure 1. The detection framework toward fine-grained LLM-generated text detection through role recognition and involvement measurement.
[…]
Illustration of the detection framework for fine-grained LLM-generated text detection, highlighting role recognition and involvement measurement.
@@ 2.2. Detection methods
Current detection methods can be broadly categorized into three types based on the features they rely on (Ghosal et al., 2023): watermarking-based detection methods, statistical outlier detection methods, and fine-tuning classifiers. (i) The watermark-based methods require embedding the signals that are invisible to humans into the AI-generated text and then detecting them based on these invisible token-level secret markers (Kirchenbauer et al., 2023; Li et al., 2024a). However, this method not only requires pre-editing that is not applicable to open-source models (Puccetti et al., 2024a) but also affects the quality of model generation due to the insertion of watermarks (Tang et al., 2024; Pang et al., [n. d.]). (ii) Statistical outlier detection methods focus on distinguishing whether a text is written by GPT based on the human features contained in the text. They adopt features ranging from shallows (entropy (Lavergne et al., 2008; Gehrmann et al., 2019), n-gram frequencies (Badaskar et al., 2008), and perplexity (Beresneva, 2016)) to deeps such as using the absolute rank (Gehrmann et al., 2019), the Log Likelihood Ratio Ranking (LRR) by complementing Log Rank (Su et al., 2023) and the model’s log probability in regions of negative curvature (DetectGPT) (Mitchell et al., 2023). (iii) Supervised fine-tuning classifiers, trained on annotated data (Jawahar et al., 2020; Bakhtin et al., 2019; Sadasivan et al., 2023), have shown effectiveness in detecting LLM-generated text across domains, such as news (Zellers et al., 2019; Kumarage et al., 2023), social media (e.g., Twitter) (Fagni et al., 2021; Puccetti et al., 2024b), and academic papers (Yang et al., 2023).
However, these classifiers often overfit to specific domains, leading to poor performance on out-of-distribution data (Uchendu et al., 2020; Chakraborty et al., 2023), and their capabilities significantly degrade when applied to unseen datasets from different domains (Li et al., 2024b). Thus, assessing the transferability of detection models is essential for their practical application across diverse datasets and domains. In our work, we evaluate detection methods—including zero-shot LLMs, supervised feature-based, and PLM-based models—on our novel tasks and dataset. Our empirical validation provides insights into improving detection models for LLM-generated content, particularly with respect to enhancing generalization and robustness in diverse, real-world scenarios.
@@ 3. METHODOLOGY
Figure 1 illustrates our proposed detection framework for fine-grained detection of LLM-generated text, encompassing two proposed novel detection tasks (§3.1), and the LLMDetect Benchmark (HNDC & DetectEval) for training and evaluation detection models (§3.2).
@@ 3.1. Detection Paradigm Definition
LLM-generated text detection currently relies on binary classification to determine whether a text is LLM-generated. Given a dataset $\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{N}$, where $x_{i}$ is the text and $y_{i}\in\left\{0,1\right\}$ indicates whether the text is LLM-generated. However, this approach focuses only on identifying LLM-generated content and is unable to distinguish more complex scenarios. To overcome this, we propose two new tasks: LLM Role Recognition (LLM-RR), a multi-class classification task, and LLM Involvement Measurement (LLM-IM), a regression task.
@@ 3.1.1. LLM Role Recognition
LLM-RR aims to identify the role of LLMs in text generation when used for assisted writing. Unlike binary detection, the label for each $x_{i}$ is $y_{i}\in{C_{1},C_{2},\dots,C_{k}}$, indicating the specific role $C_{i}$ the LLM plays. Roles include fully human-written text, LLM-generated content with minor human editing, human-led creation with LLM assistance, fully LLM-generated text, and more. The LLM-RR task can be defined as follows:
@@
where our objective is to find an optimal classifier $f(\cdot)$ that minimizes the overall average misclassification rate over the data distribution $\mathcal{D}$. The indicator function $\mathbf{1}\{f(x_{i})\neq y_{i}\}$ equals 1 when the classifier $f$ assigns an incorrect label to the input $x_{i}$, and 0 if the label is correct. The text $x_{i}$ follows a distinct distribution $\mathscr{F}_{k}$ conditioned on its category label $y_{i}$:
@@ 3.1.2. LLM Involvement Measurement
While LLM-RR offers more granularity than binary detection, it has limitations. Real-world user interactions with LLMs are complex, making it difficult to define all possible roles, and the LLM’s contribution can vary within the same role. To address these challenges, we propose LLM Involvement Measurement (LLM-IM), a regression task that quantifies the degree of LLM involvement in generated text. In this task, labels $y_{i}$ represent continuous values, known as the LLM Involvement Ratio (LIR), ranging from 0 (no LLM involvement) to 1 (entirely LLM-generated). Specifically, it is calculated as follows:
@@
3.2. LLMDetect Benchmark
To validate the effectiveness of our proposed detection paradigm, we construct LLMDetect, a benchmark specifically designed to evaluate detection models across varying levels of LLM involvement. This benchmark encompasses four distinct roles in content creation: Human-Author, LLM-Creator, LLM-Polisher, and LLM-Extender. Each of the four roles represents a distinct level of LLM participation:
@@ 3.2. LLMDetect Benchmark
Specifically, the LIR is defined as 0 for Human-Author text and 1 for LLM-Creator text, while for LLM-Polisher and LLM-Extender text, the LIR falls between 0 and 1. For LLM-Extender, the LIR value can be directly calculated using Equation 3. However, for LLM-Polisher, directly extracting $T_{LLM}$ is not feasible. Therefore, following the approach of Yang et al. (2023), we use the Jaccard distance444 The details of Jaccard distance are provided in the Appendix A. to calculate the polish ratio, which serves as the LIR for this role.
In LLMDetect, each Human-Author text is paired with three versions generated by the other LLM roles. Each text is annotated with its corresponding LLM role and associated LIR value. This dual annotation framework enables a comprehensive evaluation of both the role and extent of LLM involvement in content creation. The constructed LLMDetect benchmark comprises two key components: the Hybrid News Detection Corpus (HNDC), a diverse dataset designed for robust training and evaluation of detection methods, and DetectEval, a comprehensive evaluation suite featuring five distinct out-of-distribution settings and varying intensity levels within the same LLM role.
@@ 3.2.1. HNDC
b. LLM-Assisted News Generation
To generate LLM-generated news articles, we design distinct prompts based on the three proposed roles: (1) For LLM-Creator news, the prompt includes the title, available summary, topic category, and publication date to ensure factual reliability. (2) For LLM-Polisher news, the entire original article is provided. If the article is too long, it is segmented for polishing to avoid overly shortened outputs that may result from processing lengthy articles in one go. (3) For LLM-Extender news, we retain the first three sentences or up to one-third of the original text and instruct the LLM to generate the remaining content. To ensure high-quality generation, we employ role-playing prompts, assigning the LLM the role of a journalist. This approach leverages social role assignment, which has been shown to improve LLM performance consistently (Zheng et al., 2023).666 All prompts used for data construction are provided in the Appendix B.1. We select LLaMa3 (Dubey et al., 2024) (Meta-Llama-3-8B-Instruct) as the writing assistant LLM.
@@ 3.2.2. DetectEval
DetectEval is a comprehensive evaluation suite designed to assess the transferability and robustness of detection models, specifically focusing on five cross-context variations and two multi-intensity variations.
@@
Cross-context variations examine data diversity across five dimensions: publication time, generation prompts, LLM sources, cultural differences, and content domain, resulting in five out-of-distribution settings: cross-time, cross-prompt, cross-source, cross-cultural, and cross-domain. The first four settings, similar to HNDC, focus on LLM-generated content in the news domain. Cross-prompt and cross-source are extensions of the HNDC test dataset. (i) Cross-time: HNDC’s pre-2019 articles reduce LLM involvement, but data leakage is still possible. We scrape 2024 New York Times articles and generate LLM content for different roles using the same method. (ii) Cross-prompt: LLM-assisted writing varies by prompt. We design five prompts per role and pair them with news articles to create a diverse test set. (iii) Cross-source: HNDC used Llama-3-8B-Instruct, but real-world scenarios involve stronger models. We supplement HNDC data with content from four more powerful models: Deepseek-v2, Meta-LLaMA-3-70B-Instruct, Claude-3.5-Sonnet, and GPT-4o. (iv) Cross-cultural: To capture cultural variation in writing, we create a test set from news platforms (dai, 2017) in Germany, China, and Qatar, reflecting diverse languages and values. (v) Cross-domain: Writing styles vary across domains. We build a cross-domain dataset from thesis, story, and essay texts sourced from CUDRT (Tao et al., 2024), MGTBench (He et al., 2023), and CDB (Liang et al., 2023).
@@
Multi-intensity variations are introduced to address the fact that, even within the same LLM-assisted writing role, the level of LLM involvement can differ. We specifically construct test data with varying degrees of LLM involvement ratio for the LLM-extender and LLM-polisher roles. (i) Variable-Length Extension: For LLM-Extender role, we set up three truncation lengths to create variable-length extensions, allowing us to evaluate whether the detectors can identify the differing levels of content expansion. For the given text $x$, this process can be described as: $E^{\left(n\right)}(x),n\in\{Low,Medium,High\}$ [^7] We randomly retain part of an article’s initial sentences and ask LLMs to complete it. Low refers to retaining $[3,l/3]$ sentences, Medium retains $[l/3,2l/3]$, and High retains $[2l/3,l-3]$, where $l$ is the total number of sentences. , where $n$ denotes the truncation state of the text, and $E$ represents the process of text extension by LLMs. (ii) Multi-Staged Polish: For LLM-polisher role, we apply a multi-staged polish process, iterating the text polishing up to six times to evaluate whether the detection methods can identify the increasing levels of refinement. For the given text $x$, this process can be described as: $P_{m}\left(x\right)$, where $m$ denotes the number of polish times, and $P$ represents the polishing process by LLMs.